### v1

# Deferred Commitment Decoding for Diffusion Language Models with Confidence-Aware Sliding Windows

Abstract
Diffusion language models (DLMs) have recently emerged as a strong alternative to autoregressive models by enabling parallel text generation. To improve inference efficiency and KV-cache compatibility, prior work commonly adopts block-based diffusion, decoding tokens block by block. However, this paradigm suffers from a structural limitation that we term Boundary-Induced Context Truncation (BICT): undecoded tokens near block boundaries are forced to commit without access to nearby future context, even when such context could substantially reduce uncertainty. This limitation degrades decoding confidence and generation quality, especially for tasks requiring precise reasoning, such as mathematical problem solving and code generation. We propose Deferred Commitment Decoding (DCD), a novel, training-free decoding strategy that mitigates this issue. DCD maintains a confidence-aware sliding window over masked tokens, resolving low-uncertainty tokens early while deferring high-uncertainty tokens until sufficient contextual evidence becomes available. This design enables effective bidirectional information flow within the decoding window without sacrificing efficiency. Extensive experiments across multiple diffusion language models, benchmarks, and caching configurations show that DCD improves generation accuracy by 1.39% with comparable time on average compared to fixed block-based diffusion methods, with the most significant improvement reaching 9.0%. These results demonstrate that deferring token commitment based on uncertainty is a simple yet effective principle for improving both the quality and efficiency of diffusion language model decoding.
Deferred Commitment Decoding for Diffusion Language Models with Confidence-Aware Sliding Windows
Yingte Shu Peking University ytshu25@stu.pku.edu.cn          Yuchuan Tian Peking University tianyc@stu.pku.edu.cn          Chao Xu Peking University xuchao@cis.pku.edu.cn
Yunhe Wang Huawei Technologies Co., Ltd yunhe.wang@huawei.com          Hanting Chen Huawei Technologies Co., Ltd chenhanting@huawei.com
Figure 1: An overview of the proposed DCD algorithm.
1 Introduction
Diffusion language models (DLMs) have recently emerged as a promising alternative to autoregressive models for natural language generation. By decoding tokens in parallel rather than strictly left-to-right, DLMs relax sequential dependencies and enable more flexible generation. Recent models such as NBDiff tian2025next and LLaDA2.0 bie2025llada2 demonstrate that, at comparable scales, DLMs can match or even surpass their autoregressive counterparts on selected reasoning tasks.
A major challenge in practical DLM inference lies in compatibility with key-value (KV) caching. Vanilla DLMs decode tokens in largely unconstrained orders, which prevents effective reuse of cached attention states and leads to slow inference. To address this issue, block-based diffusion methods have been proposed arriola2025block, partitioning the sequence into blocks that are decoded sequentially while allowing parallel decoding within each block. This semi-autoregressive structure significantly improves KV-cache efficiency and has become a standard design choice in recent DLM systems.
Despite their efficiency benefits, block-based diffusion methods introduce a fundamental limitation, which we refer to as Boundary-Induced Context Truncation (BICT). Once decoding proceeds to the next block, undecoded tokens in the current block are forced to commit, even if nearby future tokens—often only a few positions away—could provide crucial disambiguating context. This issue is particularly harmful for tokens in semantically critical positions, where insufficient context leads to low-confidence decisions and error propagation. Importantly, this limitation is not caused by an incorrect decoding order, but by rigid block boundaries that assume information sufficiency at block completion.
Our core hypothesis is that decoding quality can be improved by deferring commitment on high-uncertainty tokens until sufficient contextual evidence becomes available, without abandoning the efficiency advantages of block-based decoding. Based on this insight, we propose Deferred Commitment Decoding (DCD), a training-free decoding strategy that replaces fixed block boundaries with a confidence-aware sliding window. Within this window, tokens with low uncertainty are resolved first, while high-uncertainty tokens remain masked and continue to benefit from dynamically expanding context. This mechanism enables localized bidirectional information flow while preserving compatibility with existing caching schemes.
We evaluate DCD on a diverse set of tasks, including mathematical reasoning math500; gsm8k, code generation mbpp; humaneval, and instruction following ifeval, using multiple diffusion language models nie2025large; ye2025dream; wu2025fast2 and various KV caching configurations. Across all settings, DCD consistently improves generation accuracy by 1.39% with comparable time on average compared to fixed block-based diffusion baselines, while the maximum improvement in certain configurations reaches 9.0%. These results establish DCD as a strong state-of-the-art decoding method for DLMs.
We summarize our contributions as follows:
  • We identify Boundary-Induced Context Truncation as a key structural limitation of block-based diffusion decoding, which prevents undecoded tokens from leveraging nearby future context across rigid block boundaries.
  • We propose Deferred Commitment Decoding, a simple, training-free decoding strategy that dynamically aligns the decoding order with token-level uncertainty using a sliding window.
  • We demonstrate that DCD achieves consistent accuracy improvements and decoding speedups over fixed block-based diffusion methods across models, tasks, and caching configurations.
2 Related Works
2.1 DLMs Taxonomy
There are two main lines of work that adapt diffusion techniques ho2020denoising from computer vision to natural language processing. Continuous diffusion language models li2022diffusion; gong2022diffuseq project discrete language tokens into continuous spaces and apply denoising processes to recover text outputs. In contrast, discrete diffusion language models draw inspiration from masked language modeling devlin2019bert, gradually recovering masked tokens at predefined generation slots. Compared to continuous approaches, discrete DLMs better align with the inherently discrete nature of language and can be more easily adapted from existing autoregressive models gong2024scaling; as a result, they have become the dominant paradigm in recent diffusion-based language modeling research. Unless otherwise specified, we use the term DLMs in this paper to refer to discrete diffusion language models.
Discrete DLMs typically employ one of two attention mechanisms in the Transformer architecture: semi-causal attention or full attention. Models such as BD3-LM arriola2025block and Fast-dLLMv2 wu2025fast2 adopt blockwise causal attention, where tokens attend only to their current block and preceding blocks. In contrast, models such as LLaDA nie2025large and Dream ye2025dream adopt full attention, allowing each token to condition on the entire sequence during decoding. In this work, we consider both semi-causal and full-attention DLMs to demonstrate the generality and robustness of the proposed DCD decoding algorithm across different architectural choices.
2.2 Use of Cache in DLMs
Key-value (KV) caching is a crucial optimization for Transformer-based diffusion language models, as it enables faster inference by reusing previously computed attention states. For full-attention DLMs, Fast-dLLM wu2025fast enforces a blockwise decoding order and introduces prefix and dual caching strategies, where key-value pairs are cached for prefix blocks and for both prefix and suffix blocks, respectively. dKV-cache ma2025dkv proposes a delayed KV-cache mechanism, in which tokens are cached one step after being decoded. A further variant selectively caches only the neighboring tokens of the most recently decoded positions, thereby reducing cache inconsistency and computational overhead.
Due to the bidirectional nature of full-attention DLMs, KV caching in these models is inherently approximate, necessitating periodic cache refreshes to maintain correctness. In contrast, semi-causal attention DLMs naturally support exact prefix KV caching without refresh, as later decoding steps do not alter the attention context of earlier tokens. Fast-dLLMv2 wu2025fast2 further extends this design by introducing intra-block dual caching, where both prefix and suffix “sub-block” tokens within a block are cached to improve efficiency during blockwise decoding.
Overall, these caching strategies substantially accelerate DLM inference, although they may introduce minor degradation in generation quality due to cache approximation. The proposed DCD algorithm is compatible with existing KV caching schemes in both full-attention and semi-causal DLMs, allowing it to achieve efficiency gains without sacrificing decoding accuracy.
2.3 Decoding Strategies of DLMs
Earlier works on DLMs austin2021structured; sahoo2024simple randomly unmask and remask a fixed number of tokens at each decoding step, which often yields suboptimal performance. Later approaches incorporate confidence- or entropy-based criteria, decoding tokens whose uncertainty exceeds a threshold or falls within Top-$k$ candidates. These strategies improve flexibility and parallelism but still rely on fixed decoding ranges.
[…]
3 Preliminary of DLMs Decoding
3.1 Formulations of DLMs
Discrete diffusion language models (DLMs) generate a target sequence $\mathbf{x}=(x_{1},\dots,x_{T})$ by iteratively denoising a partially masked sequence. At diffusion step $t$, the sequence $\mathbf{x}^{(t)}$ contains masked positions denoted by $\langle\text{MASK}\rangle$. The reverse denoising process is modeled as:
$p_{\theta}(\mathbf{x}^{(t-1)}\mid\mathbf{x}^{(t)})=\prod_{i\in\mathcal{M}^{(t)}}p_{\theta}(x_{i}\mid\mathbf{x}^{(t)}),$ (1)
where $\mathcal{M}^{(t)}$ denotes the set of masked positions at step $t$.
For full-attention DLMs, each masked token $x_{i}$ is predicted by conditioning on the entire partially decoded sequence $\mathbf{x}^{(t)}$. In contrast, semi-causal DLMs partition the sequence into ordered blocks $\{\mathcal{B}_{1},\dots,\mathcal{B}_{K}\}$ and restrict attention such that tokens in block $\mathcal{B}_{k}$ are conditioned only on tokens from blocks $\{\mathcal{B}_{1},\dots,\mathcal{B}_{k}\}$. Accordingly, the reverse process can be written as:
$p_{\theta}(\mathbf{x}^{(t-1)}\mid\mathbf{x}^{(t)})=\prod_{k=1}^{K}\prod_{i\in\mathcal{B}_{k}\cap\mathcal{M}^{(t)}}p_{\theta}(x_{i}\mid\mathbf{x}^{(t)}_{\leq k})$, (2)
where $\mathbf{x}^{(t)}_{\leq k}$ denotes the tokens in the first $k$ blocks.
3.2 Block-based decoding of DLMs
Decoding proceeds by selecting token values for a subset of masked positions according to the model prediction:
$x_{i}^{(t-1)}=\begin{cases}\arg\max_{v\in\mathcal{V}}p_{\theta}(v\mid\mathbf{x}^{(t)}),&i\in\mathcal{S}^{(t)},\\ x_{i}^{(t)},&\text{otherwise},\end{cases}$ (3)
where $\mathcal{V}$ is the vocabulary set and $\mathcal{S}^{(t)}=\{i\in\mathcal{M}^{(t)}\mid\bigwedge_{t}\text{Cond}_{t}(i)\}$ specifies whether position $i$ is eligible for decoding at the current step.
In block-based decoding, the eligibility condition constrains decoding positions to a fixed region, typically the current block: $\text{Cond}_{0}(i)=\bigl[i\in\mathcal{B}_{\text{cur}}\bigr]$. Full-attention DLMs may optionally adopt block-based decoding to improve KV-cache compatibility. In contrast, semi-causal DLMs must employ block-based decoding due to their blockwise attention constraints. Within a large attention block, semi-causal DLMs may further apply sub-block decoding, where decoding positions are restricted to a smaller contiguous region: $\text{Cond}_{0}(i)=\bigl[i\in\mathcal{B}_{\text{cur}_{1}:\text{cur}_{2}}\bigr]$, where $\mathcal{B}_{\text{cur}_{1}:\text{cur}_{2}}$ denotes a contiguous subrange of blocks within a larger attention block.
4 Boundary-Induced Context Truncation
The advantages of DLMs over their autoregressive counterparts lie primarily in their (semi-)bidirectional attention horizons. piskorz2025masks found that DLMs exhibit a strong contextual locality bias, in which nearby tokens contribute disproportionately to prediction confidence. Informally, this can be expressed as:
$p_{\theta}(x_{i}\mid\mathbf{x}^{(t)})\approx p_{\theta}(x_{i}\mid\mathbf{x}^{(t)}_{[i-\omega_{l}:i+\omega_{r}]})$ (4)
However, under block-based decoding, tokens after the current block are all $\langle\text{MASK}\rangle$, which contain little information and may even distract the decoding process. For tokens whose contextual locality extends beyond the current block, Equation 4 deteriorates to
$p_{\theta}(x_{i}\mid\mathbf{x}^{(t)})\approx p_{\theta}(x_{i}\mid\mathbf{x}^{(t)}_{[i-\omega_{l}:b]})$ (5)
where $b<i+\omega_{r}$ denotes the right boundary of the current block. We term this reduction in a token’s effective contextual window the Boundary-Induced Context Truncation phenomenon. Although these tokens receive insufficient context, they must be decoded before entering the next block under the block-based paradigm. Consequently, this leads to low-confidence decoding at the end of each block and ultimately degrades the generation performance of DLMs.
Will increasing the block size solve this problem? Increasing the block size partially reduces the number of blocks, but it does not break the rigidity of block boundaries. Furthermore, it may lead to the long decoding-window problem seo2025fast and weaken the effectiveness of KV caching. As additionally evidenced by previous experiments wu2025fast, setting an excessively large block size is therefore not a desirable solution.
5 Deferred Commitment Decoding
5.1 Three Rules of the DCD Algorithm
Based on the above analysis, the primary causes of BICT are rigid block boundaries and the strict left-to-right blockwise decoding order. To address this problem, we must:
  • (1)
    Identify tokens that suffer from the BICT phenomenon;
  • (2)
    Remove the restrictions imposed by rigid block boundaries;
  • (3)
    Decode these tokens at the appropriate time and under appropriate conditions.
As illustrated in Figure 1, the proposed Deferred Commitment Decoding (DCD) algorithm maintains a sliding window and defers the decoding of low-confidence tokens. It follows three rules to achieve the above goals:
Rule (1): Decoding is based on confidence criteria within the window. The prediction confidence of masked tokens serves as a key indicator for identifying BICT-affected tokens, as low confidence strongly suggests insufficient context. Similar to prior work, we set a confidence threshold $\tau_{\text{conf}}$, and only masked tokens within the decoding window whose confidence exceeds $\tau_{\text{conf}}$ are decoded. Specifically, if none of the tokens exceeds the threshold, the most confident token is decoded.
Rule (2): The decoding window slides left to right with constraints. The sliding window defines the range of tokens eligible for decoding. It abandons fixed boundaries across consecutive decoding steps; instead, it moves from left to right within the generation slot of the DLM. Formally, let $L^{(t)},R^{(t)})$ denote the left and right endpoints of the sliding window, and let $\mathbf{x}^{(t)}_{[l:r]}$ denote the generation slot at decoding step $t$. Then,
$L^{(t)}$ $=\arg\min_{i\geq l}\{i\mid x^{(t)}_{i}=\langle\text{MASK}\rangle\},$ (6)
$\begin{split}R^{(t)}&=\arg\max_{i\leq r}\{i\mid i\leq L^{(t)}+s_{\text{max}}\text{ and }\\ &\phantom{{}={}}\sum_{k=L^{(t)}}^{i-1}\left[x^{(t)}_{k}=\langle\text{MASK}\rangle\right]\leq s_{\text{init}}\}.\end{split}$ (7)
Equation 6 shows that the left endpoint of the window is anchored to the leftmost masked token. Equation 7 indicates that the sliding window expands its right boundary as much as possible, subject to two constraints: (1) the total length of the sliding window does not exceed $s_{\text{max}}$, and (2) the number of masked tokens within the window does not exceed $s_{\text{init}}$. In particular, the sliding window is initialized with length $s_{\text{init}}$ at the beginning of the generation slot.
These constraints maintain a moderate yet flexible window length, which precisely captures relevant contextual information and enables more efficient KV-cache integration.
Rule (3): Tokens are deferred until they are sufficiently confident. In block-based decoding, low-confidence tokens may be forcibly decoded to complete a block. In contrast, the proposed DCD algorithm handles low-confidence tokens more gracefully: it defers their decoding and expands the available context until the DLM becomes sufficiently confident to predict them. This approach significantly reduces low-confidence decoding events at later stages, thereby improving overall performance.
How does the DCD algorithm differ from AdaBlock? The AdaBlock method lu2025adablock employs adaptive block sizes based on delimiter semantics in the generated tokens, improving generation coherence. However, once computed, the block sizes remain fixed; as a result, AdaBlock may still suffer from BICT and thus leaves room for further improvement.
Algorithm 1 Deferred Commitment Decoding (DCD)
0: Generation slot $\mathbf{x}^{(t)}_{[l:r]}$, DLM model $p_{\theta}(\cdot\mid\mathbf{x}^{(t)})$, window parameters $s_{\text{init}},s_{\text{max}}$, cache parameters $\text{cache\_type},B^{\prime},r$, confidence threshold $\tau_{\text{conf}}$.
1: Initialize $L^{(t)},R^{(t)}=l,l+s_{\text{init}}$.
2: Initialize cache refresh countdown $cd=0$.
3:while $\mathcal{M}^{(t)}\neq\emptyset$ do
4:  Obtain eligible masked tokens $\mathcal{E}^{(t)}=[L^{(t)},R^{(t)})\cap\mathcal{M}^{(t)}$.
5:  if cache_type $\neq$ none and $cd\leq 0$ then
6:   Refresh the cache based on cache_type and Equation 8.
7:   Set $cd=B^{\prime}$.
8:  end if
9:  For all $i\in\mathcal{E}^{(t)}$, compute confidence $c_{i}=\max_{x_{i}\in\mathcal{V}}p_{\theta}(x_{i}\mid\mathbf{x}^{(t)})$.
10:  Select decoding positions $\mathcal{S}^{(t)}=\{i\mid c_{i}\geq\tau_{\text{conf}}\}\cup\arg\max_{i}\{c_{i}\}$.
11:  Update $\mathbf{x}^{(t-1)}$ with $\mathcal{S}^{(t)}$ using Equation 3.
12:  Update $L^{(t-1)},R^{(t-1)}$ with $\mathbf{x}^{(t-1)}_{[l:r]}$ using Equations 6 and 7.
13:  Update $cd=cd-|\mathcal{S}^{(t)}|$.
14:  Update t=t1t=t-1.
15:end while
16:return Final sequence $\mathbf{x}^{(t)}_{[l:r]}$.
5.2 Applying DCD to Different Types of DLMs
The generation slot of DCD is aligned with the bidirectional attention intervals of different types of DLMs.
For full-attention DLMs, there is a single generation slot consisting of the prefilled masked tokens following the prompt tokens, which represents the fixed-length output of the corresponding query. In this case, the DCD algorithm completely replaces block-based decoding.
For semi-causal DLMs, multiple generation slots arise during inference, each corresponding to a “large block” predefined by the attention structure. In this setting, the training-free DCD algorithm cannot modify the blockwise decoding pattern at the macro level, as these DLMs are trained with fixed block sizes. However, DCD can replace fixed-length sub-block decoding within each large block and outperform both standard block-based and sub-block-based decoding methods.
5.3 DCD’s Combination with KV Cache
To accelerate DLM inference, we integrate prefix and dual caching into the DCD algorithm, following Fast-dLLM wu2025fast. Inspired by dKV-Cache-Greedy ma2025dkv, the active interval without caching is slightly extended beyond the decoded tokens from the current and previous steps. Formally, it is defined as:
$\mathcal{W}^{(t)}=\left\{x^{(t)}_{i}\mid i\in[L^{(t-1)}-r,R^{(t)}+r)\right\}.$ (8)
We then define the prefix of the generation slot as the tokens preceding $\mathcal{W}^{(t)}$, and the suffix as the tokens following $\mathcal{W}^{(t)}$. The prefix cache temporarily stores the prefix, while the dual cache stores both the prefix and suffix. Additionally, to ensure a fair comparison with block-based cache refreshing, we rebuild the cache after $B^{\prime}$ masked tokens have been decoded since the previous cache refresh.
6 Experiments
Table 1: Experimental results. For each experiment, we report its overall metrics (pass@1, accuracy, etc.). We also report the total seconds for running 5 benchmarks within one line. The best result of certain model and task is bolded and second-best is underlined.
Model Cache Decoding Time Humaneval MBPP MATH500 GSM8K IFEval
(0-shot) (3-shot) (0-shot) (5-shot) (0-shot)
LLaDA-8B -Instruct Avg. Metric +1.16 Avg. Time 0.0% None Block-based 40750 43.3 39.8 40.2 78.3 57.9
Prefix Block-based 24671 43.3 39.8 38.8 76.0 56.4
Dual Block-based 18617 44.5 36.4 36.2 75.7 53.2
None DCD 40745 43.9 40.0 41.0 79.1 59.0
Prefix DCD 24803 45.7 38.2 41.2 78.5 57.1
Dual DCD 18501 44.5 37.2 39.0 79.2 53.6
dKV-Cache-Greedy - 15.37 20.4 27.0 68.23 -
Dual AdaBlock 23680 45.1 36.2 36.6 78.4 55.8
Dream-v0- Instruct-7B Avg. Metric +2.63 Avg. Time -2.2% None Block-based 23685 54.3 55.0 44.8 76.6 50.5
Prefix Block-based 15420 56.7 53.6 43.4 77.6 51.8
Dual Block-based 9273 56.7 52.8 44.4 74.8 47.7
None DCD 23449 53.7 56.8 43.8 78.2 55.6
Prefix DCD 15044 58.5 57.4 43.4 78.6 56.4
Dual DCD 9284 59.8 58.8 45.2 77.3 56.7
Dream-v0- base-7B Avg. Metric +0.77 Avg. Time -9.2% None Block-based 25594 48.2 13.8 12.0 75.5 -
Prefix Block-based 14600 57.3 13.6 12.6 74.5 -
Dual Block-based 10189 57.3 13.4 13.2 73.8 -
None DCD 22714 50.6 17.0 12.8 76.0 -
Prefix DCD 13283 53.0 16.0 12.8 74.4 -
Dual DCD 9406 56.1 13.2 13.2 74.7 -
Dual AdaBlock 42141 53.0 14.4 13.0 76.0 -
Fast-dLLM -v2-7B Avg. Metric +0.62 Avg. Time -8.3% None Block-based 9993 56.7 48.4 50.8 74.5 62.5
None Sub-block-based 11228 61.0 50.2 54.6 77.6 62.8
Dual Sub-block-based 11379 57.9 46.0 52.4 76.0 60.3
None DCD 10116 62.8 48.6 53.4 77.9 64.0
Dual DCD 10498 59.1 49.0 51.6 77.8 60.8
Figure 2: Confidence distributions of LLaDA-8B-Instruct, Dream-v0-Base-7B, and Fast-dLLM-v2-7B on GSM8K with DCD and (sub-)block-based decoding using dual cache. We use a log10 scale for clarity. The DCD algorithm yields fewer low-confidence decoding steps across these models.
Figure 3: Ablation studies on LLaDA-8B-Instruct on the MATH500 task with dual-cache DCD. We vary $s_{\text{max}}$, $s_{\text{init}}$, and $\tau_{\text{conf}}$, and evaluate task accuracy and decoding time.
6.1 Experimental Setup
Models.
To validate the effectiveness of the DCD algorithm, we evaluate four pretrained DLMs: LLaDA-8B-Instruct nie2025large, Dream-v0-Instruct-7B, Dream-v0-Base-7B ye2025dream, and Fast-dLLM-v2-7B wu2025fast2. The first three models adopt full attention, while the last model uses semi-causal attention.
Benchmarks.
For each model, we evaluate coding benchmarks including HumanEval humaneval and MBPP mbpp, mathematical reasoning benchmarks including MATH500 math500 and GSM8K gsm8k, and the instruction-following benchmark IFEval ifeval. The Dream-v0-Base-7B model is excluded from IFEval because it is not instruction-aligned. We do not evaluate multiple-choice QA benchmarks mmlu; gpqa, as they primarily measure token-level log-probabilities rather than decoding quality.
Cache Configurations and Baselines.
For full-attention DLMs, the parallel block-based decoding of Fast-dLLM wu2025fast serves as the baseline under three cache configurations: no cache, prefix cache, and dual cache. For the semi-causal DLM Fast-dLLM-v2-7B wu2025fast2, sub-block-based decoding with no cache and with dual cache within each large block is used as the primary baseline, while vanilla block-based decoding without sub-block structures is used as an additional baseline. For each model, benchmark, and cache configuration, DCD is compared against these baselines as well as other training-free DLM decoding strategies such as dKV-Cache ma2025dkv and AdaBlock-dLLM lu2025adablock, when available.
Hyperparameters.
For full-attention DLMs, we set the generation slot length to $L=512$, window parameters to $s_{\text{init}}=16$ and $s_{\text{max}}=128$, cache parameters to $B^{\prime}=32$ and $r=2$, and the confidence threshold to $\tau_{\text{conf}}=0.9$. For the semi-causal DLM, we set $s_{\text{init}}=8$ and $s_{\text{max}}=\infty$, while keeping all other parameters unchanged. All baselines use a block size of $B=32$ across all four models, with a sub-block size of $b=8$ for Fast-dLLM-v2-7B wu2025fast2. The decoding temperature is set to 0 and the batch size is set to 1 for all experiments. Additional details are provided in Appendix C.
6.2 Main Results Analysis
Table 1 reports the main experimental results across multiple diffusion language models, tasks, and KV-cache configurations. Overall, Deferred Commitment Decoding (DCD) consistently outperforms block-based and sub-block-based decoding across most settings, demonstrating strong robustness across tasks and model architectures. On average, DCD improves evaluation metrics by +1.39% while reducing decoding time by 4.4% compared to block-based (or sub-block-based) baselines across all models. The average improvement for each model against baseline is highlighted in colored text in Table 1.
These gains are observed across diverse tasks, including mathematical reasoning, code generation, and instruction following, and across both full-attention and semi-causal DLMs. Among all models, Dream-v0-Instruct-7B benefits the most from DCD, achieving the largest average improvement of 2.53 points. Running the MBPP benchmark with Dream-v0-Instruct-7B and dual cache yields the most significant improvement, with a 9.0% increase in pass@1. In contrast, Fast-dLLM-v2-7B shows the smallest improvement of 0.62 points. This outcome is expected, as DCD currently operates only at the sub-block level for semi-causal models and does not modify the macro-level block-based decoding paradigm imposed by their attention structure. Nevertheless, DCD still consistently outperforms sub-block-based decoding, which aligns with our theoretical analysis.
Compared with AdaBlock lu2025adablock, DCD achieves better performance (average metric improvements of +0.28% for LLaDA-8B-Instruct and +1.20% for Dream-v0-Base-7B) and substantial speedups (average decoding time reductions of 19% and 71%, respectively), demonstrating the effectiveness of the deferred commitment mechanism. Compared with dKV-Cache-Greedy ma2025dkv, DCD substantially outperforms it in terms of accuracy, despite employing a similar KV-cache strategy.
We note that in a small number of cases, DCD performs slightly worse than block-based decoding. We attribute these regressions to the inherent stochasticity of training-free decoding and the intrinsic difficulty of certain tokens, which may remain ambiguous even with extended context.
6.3 Evidence of BICT Mitigation
Figure 2 provides direct evidence that DCD mitigates Boundary-Induced Context Truncation. We visualize the distribution of decoding confidence on the GSM8K benchmark for LLaDA-8B-Instruct, Dream-v0-Base-7B, and Fast-dLLM-v2-7B. GSM8K is selected because it is the largest benchmark and exhibits the most stable improvements under DCD.
Across all models, DCD substantially reduces the frequency of extremely low-confidence decoding steps compared to block-based or sub-block-based decoding. Such low-confidence events directly reflect the BICT phenomenon that DCD is designed to address. This reduction provides a clear explanation for the observed accuracy improvements, particularly on reasoning-intensive tasks.
6.4 Ablation Studies
Figure 3 presents ablation studies on LLaDA-8B-Instruct evaluated on MATH500 with dual cache enabled, analyzing the impact of three key hyperparameters in Deferred Commitment Decoding (DCD): the maximum window size $s_{\max}$, the initial window size $s_{\text{init}}$, and the confidence threshold $\tau_{\text{conf}}$.
Effect of $s_{\max}$.
The maximum window size determines the upper bound of contextual expansion. When $s_{\max}=32$, the window is constrained to the baseline block size, limiting DCD’s ability to mitigate BICT. When $s_{\max}=512$ (equivalent to removing the upper bound given $L=512$), accuracy degrades due to diluted contextual relevance. Decoding time does not exhibit a consistent monotonic trend with respect to $s_{\max}$, as the window rarely expands to its maximum in practice.
Effect of $s_{\text{init}}$.
The initial window size controls early decoding behavior. Setting $s_{\text{init}}=8$ reduces the available context and may degrade performance, while $s_{\text{init}}=32$ may lead to excessively long windows, introducing premature commitments near the right boundary. The weak negative correlation between $s_{\text{init}}$ and decoding time may result from increased parallelism enabled by larger windows.
Effect of $\tau_{\text{conf}}$.
The confidence threshold regulates how aggressively tokens are deferred. Accuracy improves as $\tau_{\text{conf}}$ increases from lower values, but degrades when $\tau_{\text{conf}}=1$ (i.e., top-1 confidence decoding), as this setting becomes fully deterministic and loses flexibility. In contrast to window parameters, $\tau_{\text{conf}}$ exhibits a clearer positive correlation with decoding time.
Overall, accuracy exhibits a clear unimodal trend as $s_{\max}$, $s_{\text{init}}$, and $\tau_{\text{conf}}$ increase in this setting. Based on these ablation results, we select appropriate hyperparameters and apply them consistently across all experiments.
7 Conclusion
In this work, we investigate a fundamental limitation of block-based diffusion decoding for language models, which we formalize as Boundary-Induced Context Truncation. We identify suboptimal token commitment, whereby tokens that would otherwise benefit from nearby future context are forced to commit at block boundaries, leading to low-confidence predictions and degraded generation quality. To address this issue, we propose Deferred Commitment Decoding, a simple, training-free decoding strategy that replaces fixed block boundaries with a confidence-aware sliding window. By deferring uncertain tokens until sufficient context becomes available, DCD enables more effective utilization of local bidirectional context without sacrificing KV-cache compatibility. Extensive experiments across multiple DLMs, benchmarks, and caching configurations demonstrate that DCD consistently improves generation accuracy by 1.39% with comparable time on average compared to fixed block-based baselines, with the maximum improvement reaching 9.0%. These results demonstrate the superiority of DCD for DLM inference.
Limitations
A major limitation of the proposed DCD method arises in semi-causal DLMs, where we only modify the sub-block structure and consequently observe smaller improvements than in full-attention DLMs. We hope that future work will adapt semi-causal DLM architectures to better accommodate the DCD mechanism and achieve stronger performance across a wider range of benchmarks.
Appendix A Experimental Environment
Hardware.
For comparison experiments between block-based decoding and DCD, we allocate one NVIDIA A100 80G GPU and eight Intel(R) Xeon(R) Platinum 8358 CPUs for each experiment. For AdaBlock experiments, we allocate one NVIDIA A800 GPU and sixteen Intel(R) Xeon(R) Platinum 8378A CPUs for each experiment.
Software.
We use Ubuntu Linux, CUDA 12.0, Python 3.10, and lm-eval-harness 0.4.8 for all experiments.
Appendix B A Case Study of the BICT Phenomenon
Figure 4: MBPP Task #129 solved by block-based decoding and the DCD algorithm using Dream-v0-Instruct-7B.
Figure 4 illustrates the BICT phenomenon in block-based decoding. This example corresponds to the 129th test case of the MBPP benchmark, generated by Dream-v0-Instruct-7B without any caching. The prompt for this task is: “Write a function to extract elements that occur singly in the given tuple list.” The test cases are:
1assert extract_singly([(3, 4, 5), (4, 5, 7), (1, 4)]) == [3, 4, 5, 7, 1]
2assert extract_singly([(1, 2, 3), (4, 2, 3), (7, 8)]) == [1, 2, 3, 4, 7, 8]
3assert extract_singly([(7, 8, 9), (10, 11, 12), (10, 11)]) == [7, 8, 9, 10, 11, 12]
The incorrect code generated by the block-based method is:
1def extract_singly(test_tup):
2 res = []
3 for tup in test_tup:
4 for i in tup:
5 if i not in tup:
6 res.append(i)
7 return res
The error occurs at decoding step 21, as illustrated in the left part of Figure 4. The $\langle\text{MASK}\rangle$ at a critical position (the rightmost token of line 5) is adjacent to the first block boundary and suffers from truncated context. As a result, it is incorrectly decoded as “tup” with low confidence.
The correct code generated by the DCD method is:
1def extract_singly(test_tup):
2 res = []
3 for tup in test_tup:
4 for num in tup:
5 if num not in res:
6 res.append(num)
7 return res
The critical decoding step 23 is illustrated in the right part of Figure 4. The DCD algorithm successfully resolves this issue by deferring the decoding of the critical $\langle\text{MASK}\rangle$ until the sliding window incorporates sufficient contextual information, such as “res.append”. Consequently, the model correctly decodes the second occurrence of “res” and successfully passes all test cases.
Appendix C Details about Main Experiments
All evaluation scores are generated using lm-eval-harness 0.4.8 and the code-cleaning suite in the Fast-dLLM codebase. Specifically:
  • For HumanEval, we use 0-shot pass@1 as the evaluation metric. For code cleaning, we concatenate the prompt and the generated output, remove possible “‘python … “‘ blocks, and extract the function based on Python syntax trees. No prompt engineering is applied in this setting.
  • For MBPP, we use 3-shot pass@1 as the evaluation metric. The code-cleaning logic is the same as that for HumanEval. Prompt engineering and the few-shot mechanism are automatically handled by the lm-eval-harness library.
  • For MATH500, we use 0-shot accuracy as the evaluation metric, with simple chain-of-thought reasoning prompts as the default. The cleaning logic extracts the boxed answer and simplifies the mathematical expression.
[…]
    Respond in the following format:
    <reasoning>
    Your reasoning here
    </reasoning>
    <answer>
    boxed{...}
    </answer>
  • For GSM8K, we use 5-shot accuracy as the evaluation metric, corresponding to the “exact_match,flexible-extract” value reported in the lm-eval-harness result files. All other settings follow the default configuration.
  • For IFEval, we use 0-shot accuracy as the evaluation metric, corresponding to the “prompt_level_strict_acc,none” value reported in the lm-eval-harness result files. All other settings follow the default configuration.
Notably, we do not use any stop words (e.g., “[DONE]” in the default MBPP configuration) in any experiments. This choice may lead to degraded performance for Dream-v0-Base-7B on these tasks, as unaligned models often struggle to terminate generation appropriately and may produce extraneous content after completing the task. However, this setting is applied consistently across all experiments in this paper, ensuring fair comparisons.
Appendix D Time Consumption of Each Experiment
To better understand the DCD method, we record the time consumption for each benchmark experiment. According to Table 2, DCD completes benchmarks in comparable and even slightly less time than the baseline method, demonstrating its efficiency against traditional methods.
Table 2: Detailed time consumption for each experiment. (unit: seconds)
Model Cache Decoding Humaneval MBPP MATH500 GSM8K IFEval
(0-shot) (3-shot) (0-shot) (5-shot) (0-shot)
LLaDA-8B -Instruct None Block-based 2128 7333 5958 18144 7187
Prefix Block-based 1536 3778 4551 8584 6222
Dual Block-based 1289 2934 3594 6356 4444
None DCD 2027 7467 5702 18167 7382
Prefix DCD 1565 3824 4478 8552 6384
Dual DCD 1252 2877 3620 6266 4486
dKV-Cache-Greedy Not Available
Dual AdaBlock 1531 3934 3983 9350 4882
Dream-v0- Instruct-7B None Block-based 1038 1688 5106 10186 5667
Prefix Block-based 821 1022 3912 4922 4743
Dual Block-based 477 587 2587 2807 2815
None DCD 986 1588 5000 10299 5576
Prefix DCD 783 948 3790 4844 4679
Dual DCD 470 586 2532 2879 2817
Dream-v0- base-7B None Block-based 1534 6777 5602 11681 -
Prefix Block-based 1177 3963 3943 5517 -
Dual Block-based 717 2684 3594 3194 -
None DCD 1209 6987 5231 9287 -
Prefix DCD 1066 4053 3729 4435 -
Dual DCD 658 2415 3620 2713 -
Dual AdaBlock 2243 7105 6432 26361 -
Fast-dLLM -v2-7B None Block-based 561 1487 2168 2831 2946
None Sub-block-based 615 1566 2650 3285 3112
Dual Sub-block-based 670 1667 2428 3451 3163
None DCD 561 1519 2253 2793 2990
Dual DCD 598 1608 2339 2895 3058